OpenStudy (anonymous):

Find the vertex of the function y = -x2 + x Question 8 options: 1) (1/2, 1/4) 2) (1/4, 1/2) 3) (1/2, 1/2) 4) (1/4, 1/4)

OpenStudy (mathmale):

Please write "x squared" as x^2 for clarity, or, using Equation Editor, write \[x^2.\] The relevant equation of a parabola here is the quadratic equation y=ax^2+bx+c. compare this general equation to the given y=-x^2+1x+0. What are the values of a, b and c? The x-coordinate of the vertex is x=-b/(2a). Once you have this x value, substitute it into the given equation to find the y value at the vertex. Finally, write the vertex in the form \[(x _{v},y _{v})\]